OpenStudy (dls):

A solid conducting sphere of Radius R and total charge q rotates about its diametric axis with constant Angular speed omega,Magnetic moment of the sphere=?

OpenStudy (dls):

@Mashy

OpenStudy (anonymous):

do u know the expression for magnetic moment? :P

OpenStudy (dls):

\[\Huge M=N I A\]

OpenStudy (dls):

!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (dls):

=IA

OpenStudy (anonymous):

ok.. good boy :P.. Well.. here.. its gonna be a lil complicated ..u need to do an integral!

OpenStudy (dls):

integration wagera karna hai shayad jo nahi aata

OpenStudy (dls):

^^

OpenStudy (anonymous):

lemme thinkaa!

OpenStudy (anonymous):

u know how to find the current right?! imagine.. a small cross sectional ring on the sphere.. let dq be the charge on that ring.. then the current di = dq/(2*3.14/omega)

OpenStudy (anonymous):

ahh!.. that ain't gonna work!

OpenStudy (dls):

:/ Since the question says "Conducting sphere" what does that mean?charge will only reside on the outer surface?If yes,why so?what will be the difference in solving the question if it was not conducting?

OpenStudy (anonymous):

the conducting spehre makes sure the charge is uniformly distributed over the outer surface.. supposedly making it simpler.. but i seriously don't understand how to!..

OpenStudy (anonymous):

do u have the solution?! is the answer \[M = \frac{wR^2q}{8}\]

OpenStudy (dls):

umm no..

OpenStudy (anonymous):

is it \[M = \frac{\pi^4wR^2q}{}\]

OpenStudy (anonymous):

divided by 2 i mean!

OpenStudy (dls):

umm no

OpenStudy (anonymous):

whats the answer?! :P

OpenStudy (dls):

\[\Huge M=\frac{q \omega R^2 }{5}\]

OpenStudy (anonymous):

lemme try again!

OpenStudy (dls):

okay :|

OpenStudy (anonymous):

i dunno.. this would involve a double integral i think.. seriously which book are you referring?!

OpenStudy (dls):

what on earth is this? :| and dunno..and DC pande's question I guess..similar question in HC verma

OpenStudy (experimentx):

or \[ \int_0^{\pi } \pi (R \sin \theta)^2 \cdot \sigma ( R d\theta \cdot 2 \pi R \sin \theta ) \cdot \frac{1}{T } \] where \[ T = \frac{2 \pi }{\omega } \]

OpenStudy (experimentx):

ans \( \sigma = Q/4\pi R^2 \)

OpenStudy (experimentx):

*and

OpenStudy (dls):

i already posted the answer!

OpenStudy (experimentx):

the integral evaluates to \[ \frac 1 3 Q \omega R^2\] http://www.wolframalpha.com/input/?i=Integrate%5B%5C%5BPi%5D+%28R+Sin%5Bx%5D%29%5E2+Q%2F%284+%5C%5BPi%5D+R%5E2%29+%28R%5E2+Sin%5B++++++x%5D+%5C%5BOmega%5D%29%2C+%7Bx%2C+0%2C+Pi%7D%5D check it out here. http://lynx.uio.no/trine/fys3510/oppg_3_20_eng.pdf \[ \Huge M=\frac{Q \omega R^2 }{5} \] is magnetic moment of a uniformly charged (non conducting) sphere. For conducting sphere, the charges only lie on the surface. possibly you made error on your question or your answer sheet is wrong. or I misunderstood the above literature on that paper.